# Model Selection for Off-policy Evaluation: New Algorithms and Experimental Protocol

### v2

@@ C.3 Proof of Theorem 3
$J(\pi)-\operatorname{\mathbb{E}}_{d_{0}}\left[\widehat{Q}(s,\pi)\right]$ $\leq\frac{1}{1-\gamma}\sqrt{\mathcal{C}^{\pi}\cdot\operatorname{\mathbb{E}}_{\mu}\left[\left(\left[\mathcal{T}^{\pi}\widehat{Q}\right](s,a)-\widehat{Q}(s,a)\right)^{2}\right]}$ (80)
$\leq\frac{1}{1-\gamma}\sqrt{8\cdot\mathcal{C}^{\pi}\cdot\left(\varepsilon_{\mathsf{obj}}+\varepsilon_{\mathsf{reg}}^{2}\right)},$ (81)
$=\frac{V_{\max}}{1-\gamma}\sqrt{\frac{152\cdot\mathcal{C}^{\pi}\cdot\log\left(\frac{2m}{\delta}\right)}{n}},$ (82)
@@
$\mathbb{V}_{\mu}\left[\left(g(s,a)-Q(s,a)\right)^{2}\right]$ $\leq\operatorname{\mathbb{E}}_{\mu}\left[\left(g(s,a)-Q(s,a)\right)^{4}\right]$ (83)
$\leq V_{\max}^{2}\cdot\operatorname{\mathbb{E}}_{\mu}\left[\left(g(s,a)-Q(s,a)\right)^{2}\right].$ (84)
@@
$\left\lvert\operatorname{\mathbb{E}}_{\mu}\left[\left(g(s,a)-Q(s,a)\right)^{2}\right]-\operatorname{\mathbb{E}}_{\mathcal{D}}\left[\left(g(s,a)-Q(s,a)\right)^{2}\right]\right\rvert$ (85)
$\leq\sqrt{\frac{4\mathbb{V}_{\mu}\left[\left(g(s,a)-Q(s,a)\right)^{2}\right]\log\left(\frac{2m}{\delta}\right)}{n}}+\frac{V_{\max}^{2}\log\left(\frac{2m}{\delta}\right)}{n}$ (86)
$\leq\sqrt{\frac{4V_{\max}^{2}\mathbb{E}_{\mu}\left[\left(g(s,a)-Q(s,a)\right)^{2}\right]\log\left(\frac{2m}{\delta}\right)}{n}}+\frac{V_{\max}^{2}\log\left(\frac{2m}{\delta}\right)}{n}$ (87)
$\leq\frac{\mathbb{E}_{\mu}\left[\left(g(s,a)-Q(s,a)\right)^{2}\right]}{2}+\frac{3V_{\max}^{2}\log\left(\frac{2m}{\delta}\right)}{n}.$ (88)
@@
$\frac{1}{2}\cdot\operatorname{\mathbb{E}}_{\mu}\left[\left(g(s,a)-Q(s,a)\right)^{2}\right]$ $\leq\operatorname{\mathbb{E}}_{\mathcal{D}}\left[\left(g(s,a)-Q(s,a)\right)^{2}\right]+\frac{3V_{\max}^{2}\log\left(\frac{2m}{\delta}\right)}{n},$ (89)
$\operatorname{\mathbb{E}}_{\mathcal{D}}\left[\left(g(s,a)-Q(s,a)\right)^{2}\right]$ $\leq\frac{3}{2}\cdot\operatorname{\mathbb{E}}_{\mu}\left[\left(g(s,a)-Q(s,a)\right)^{2}\right]+\frac{3V_{\max}^{2}\log\left(\frac{2m}{\delta}\right)}{n}.$ (90)
@@ C.4 Proof of Theorem 4
$J(\pi)-\operatorname{\mathbb{E}}_{d_{0}}\left[\widehat{Q}(s,\pi)\right]$ $=\operatorname{\mathbb{E}}_{d_{0},\pi}\left[Q^{\pi}(s,a)-\widehat{Q}(s,a)\right]$ (91)
$=\frac{1}{1-\gamma}\operatorname{\mathbb{E}}_{d^{\pi}}\left[Q^{\pi}(s,a)-\gamma Q^{\pi}(s^{\prime},\pi)-\widehat{Q}(s,a)-\gamma\widehat{Q}(s^{\prime},\pi)\right]$ (92)
$=\frac{1}{1-\gamma}\operatorname{\mathbb{E}}_{d^{\pi}}\left[Q^{\pi}(s,a)-\left[\mathcal{T}^{\pi}Q^{\pi}\right](s,a)-\widehat{Q}(s,a)+\left[\mathcal{T}^{\pi}\widehat{Q}\right](s,a)\right]$ (93)
$=\frac{1}{1-\gamma}\operatorname{\mathbb{E}}_{d^{\pi}}\left[\left[\mathcal{T}^{\pi}\widehat{Q}\right](s,a)-\widehat{Q}(s,a)\right]$ (94)
$\leq\frac{\mathcal{C}_{\infty}^{\pi}}{1-\gamma}\cdot\operatorname{\mathbb{E}}_{\mu}\left[\left\lvert\left[\mathcal{T}^{\pi}\widehat{Q}\right](s,a)-\widehat{Q}(s,a)\right\rvert\right]$ (95)
$\leq\max_{g\in\mathcal{G}_{\widehat{Q}}}\operatorname{\mathbb{E}}_{\mu}\left[\textrm{sgn}\left(\widehat{Q}(s,a)-g(s,a)\right)\left(\widehat{Q}(s,a)-r-\gamma\widehat{Q}(s^{\prime},\pi)\right)\right]$ (96)
@@
${\textrm{sgn}\left(Q(s,a)-g(s,a)\right)\left(Q(s,a)-r-\gamma Q(s^{\prime},\pi)\right)}\in\left[-V_{\max},V_{\max}\right],$ (97)
@@
$\left\lvert\operatorname{\mathbb{E}}_{\mu}\left[\textrm{sgn}\left(\widehat{Q}(s,a)-g(s,a)\right)\left(\widehat{Q}(s,a)-r-\gamma\widehat{Q}(s^{\prime},\pi)\right)\right]-\operatorname{\mathbb{E}}_{\mathcal{D}}\left[\textrm{sgn}\left(\widehat{Q}(s,a)-g(s,a)\right)\left(\widehat{Q}(s,a)-r-\gamma\widehat{Q}(s^{\prime},\pi)\right)\right]\right\rvert$ (98)
$\qquad\qquad\qquad\leq 2V_{\max}\sqrt{\frac{\log\left(\frac{2m}{\delta}\right)}{n}}\vcentcolon={}\varepsilon_{\mathsf{obj}}.$ (99)
@@
$J(\pi)-\operatorname{\mathbb{E}}_{d_{0}}\left[\widehat{Q}(s,\pi)\right]$ $\leq\max_{g\in\mathcal{G}_{\widehat{Q}}}{}\operatorname{\mathbb{E}}_{\mathcal{D}}\left[\textrm{sgn}\left(\widehat{Q}(s,a)-g(s,a)\right)\left(\widehat{Q}(s,a)-r-\gamma\widehat{Q}(s^{\prime},\pi)\right)\right]+\varepsilon_{\mathsf{obj}}$ (100)
$\leq\max_{g\in\mathcal{G}_{Q^{\pi}}}{}\operatorname{\mathbb{E}}_{\mathcal{D}}\left[\textrm{sgn}\left(Q^{\pi}(s,a)-g(s,a)\right)\left(Q^{\pi}(s,a)-r-\gamma Q^{\pi}(s^{\prime},\pi)\right)\right]+\varepsilon_{\mathsf{obj}}$ (101)
$\leq\max_{g\in\mathcal{G}_{Q^{\pi}}}{}\operatorname{\mathbb{E}}_{\mu}\left[\textrm{sgn}\left(Q^{\pi}(s,a)-g(s,a)\right)\left(Q^{\pi}(s,a)-r-\gamma Q^{\pi}(s^{\prime},\pi)\right)\right]+2\cdot\varepsilon_{\mathsf{obj}}$ (102)
$=\max_{g\in\mathcal{G}_{Q^{\pi}}}{}\operatorname{\mathbb{E}}_{\mu}\left[\textrm{sgn}\left(Q^{\pi}(s,a)-g(s,a)\right)\left(Q^{\pi}(s,a)-\left[\mathcal{T}^{\pi}Q^{\pi}\right](s,a)\right)\right]+2\cdot\varepsilon_{\mathsf{obj}}$ (103)
$=\max_{g\in\mathcal{G}_{Q^{\pi}}}{}\operatorname{\mathbb{E}}_{\mu}\left[\textrm{sgn}\left(Q^{\pi}(s,a)-g(s,a)\right)\left(Q^{\pi}(s,a)-Q^{\pi}(s,a)\right)\right]+2\cdot\varepsilon_{\mathsf{obj}}$ (104)
$=2\cdot\varepsilon_{\mathsf{obj}},$ (105)
@@
$\widehat{Q}=\operatorname*{arg\,min}_{Q\in\mathcal{Q}}\max_{g\in\mathcal{G}_{Q}}\operatorname{\mathbb{E}}_{\mathcal{D}}\left[\textrm{sgn}\left(Q(s,a)-g(s,a)\right)\left(Q(s,a)-r-Q(s^{\prime},\pi)\right)\right].$ (106)
@@
$J(\pi)-\operatorname{\mathbb{E}}_{d_{0}}\left[\widehat{Q}(s,\pi)\right]\leq 2\cdot\mathcal{C}_{\infty}^{\pi}\cdot\varepsilon_{\mathsf{obj}}.$ (107)
@@ C.5 Further Discussion on Comparison to BVFT and the Coverage Assumptions
As mentioned above, the coverage condition of LSTD-Tournament is inherited from its “base” algorithm, i.e., the $\sigma_{\min}(A)$ term in LSTDQ. Despite its wide use in the analyses of LSTDQ algorithms [28, 3, 46], we discover several problems with this definition that warrants further investigation. First, the quantity is not invariant to reparameterization, in the sense that if we scale the linear features $\phi$ with a constant (and scale $\theta$ accordingly), $\sigma_{\min}(A)$ will also change despite that the estimation problem essentially remains the same. Second, if $\phi$ has linearly dependent coordinates, $\sigma_{\min}(A)$ can be 00 while this says nothing about the quality of data (which could have perfect coverage); this is usually not a concern in LSTDQ analysis since we can always pre-process the features, but for its application in LSTD-Tournament the lack of linear independence can happen if we have nearly identical candidate functions $Q_{i}\approx Q_{j}$. Last but not least, we lack good understanding of when $\sigma_{\min}(A)$ is well-behaved outside very special cases; for example, it is well-known that fully on-policy data (in the sense that $\mu$ is invariant to the transition dynamics under $\pi$) implies lower-bounded $\sigma_{\min}(A)$. However, as soon as $\mu$ is off-policy, there lacks general characterization of what kind of $\mu$ helps lower-bound $\sigma_{\min}(A)$. For example, when concentrability coefficient $\mathcal{C}_{\infty}^{\pi}$ is well bounded, it is unclear whether that implies lower-bounded $\sigma_{\min}(A)$.
@@ D.2 Experiment Settings
In the above setup, the behavior and the target policies all stem from the same DDPG training procedure. While these policies still have significant differences (see Figure 1L), the distribution shift is relatively mild. For the data coverage experiments (Section 6.3), we prepare a different set of behavior policies that intentionally offer poor coverage: these policies, denoted as $\pi_{i}^{\textrm{poor}}$, are obtained by running DDPG with a different neural architecture (than the one used for generating $\pi_{0:14}$) in a different environment of $g=-60,\sigma=100$. We also provide the parallel of our main experiments in Figures 2 and 3 under these behavior policies with poor coverage in Appendix E.1.
@@
Gravity $\mathbf{g}$ Noise Level $\sigma$ Groundtruth Model $M^{\star}$ and Behavior Policy $\pi_{b}$ Target Policies $\Pi$

### v3

@@ C.3 Proof of Theorem 3
$J(\pi)-\operatorname{\mathbb{E}}_{d_{0}}\left[\widehat{Q}(s,\pi)\right]$ $\leq\frac{1}{1-\gamma}\sqrt{\mathcal{C}^{\pi}\cdot\operatorname{\mathbb{E}}_{\mu}\left[\left(\left[\mathcal{T}^{\pi}\widehat{Q}\right](s,a)-\widehat{Q}(s,a)\right)^{2}\right]}$
$\leq\frac{1}{1-\gamma}\sqrt{8\cdot\mathcal{C}^{\pi}\cdot\left(\varepsilon_{\mathsf{obj}}+\varepsilon_{\mathsf{reg}}^{2}\right)},$
$=\frac{V_{\max}}{1-\gamma}\sqrt{\frac{152\cdot\mathcal{C}^{\pi}\cdot\log\left(\frac{2m}{\delta}\right)}{n}},$
@@
$\mathbb{V}_{\mu}\left[\left(g(s,a)-Q(s,a)\right)^{2}\right]$ $\leq\operatorname{\mathbb{E}}_{\mu}\left[\left(g(s,a)-Q(s,a)\right)^{4}\right]$
$\leq V_{\max}^{2}\cdot\operatorname{\mathbb{E}}_{\mu}\left[\left(g(s,a)-Q(s,a)\right)^{2}\right].$
@@
$\left\lvert\operatorname{\mathbb{E}}_{\mu}\left[\left(g(s,a)-Q(s,a)\right)^{2}\right]-\operatorname{\mathbb{E}}_{\mathcal{D}}\left[\left(g(s,a)-Q(s,a)\right)^{2}\right]\right\rvert$
$\leq\sqrt{\frac{4\mathbb{V}_{\mu}\left[\left(g(s,a)-Q(s,a)\right)^{2}\right]\log\left(\frac{2m}{\delta}\right)}{n}}+\frac{V_{\max}^{2}\log\left(\frac{2m}{\delta}\right)}{n}$
$\leq\sqrt{\frac{4V_{\max}^{2}\mathbb{E}_{\mu}\left[\left(g(s,a)-Q(s,a)\right)^{2}\right]\log\left(\frac{2m}{\delta}\right)}{n}}+\frac{V_{\max}^{2}\log\left(\frac{2m}{\delta}\right)}{n}$
$\leq\frac{\mathbb{E}_{\mu}\left[\left(g(s,a)-Q(s,a)\right)^{2}\right]}{2}+\frac{3V_{\max}^{2}\log\left(\frac{2m}{\delta}\right)}{n}.$
@@
$\frac{1}{2}\cdot\operatorname{\mathbb{E}}_{\mu}\left[\left(g(s,a)-Q(s,a)\right)^{2}\right]$ $\leq\operatorname{\mathbb{E}}_{\mathcal{D}}\left[\left(g(s,a)-Q(s,a)\right)^{2}\right]+\frac{3V_{\max}^{2}\log\left(\frac{2m}{\delta}\right)}{n},$
$\operatorname{\mathbb{E}}_{\mathcal{D}}\left[\left(g(s,a)-Q(s,a)\right)^{2}\right]$ $\leq\frac{3}{2}\cdot\operatorname{\mathbb{E}}_{\mu}\left[\left(g(s,a)-Q(s,a)\right)^{2}\right]+\frac{3V_{\max}^{2}\log\left(\frac{2m}{\delta}\right)}{n}.$
@@ C.4 Proof of Theorem 4
$J(\pi)-\operatorname{\mathbb{E}}_{d_{0}}\left[\widehat{Q}(s,\pi)\right]$ $=\operatorname{\mathbb{E}}_{d_{0},\pi}\left[Q^{\pi}(s,a)-\widehat{Q}(s,a)\right]$
$=\frac{1}{1-\gamma}\operatorname{\mathbb{E}}_{d^{\pi}}\left[Q^{\pi}(s,a)-\gamma Q^{\pi}(s^{\prime},\pi)-\widehat{Q}(s,a)-\gamma\widehat{Q}(s^{\prime},\pi)\right]$
$=\frac{1}{1-\gamma}\operatorname{\mathbb{E}}_{d^{\pi}}\left[Q^{\pi}(s,a)-\left[\mathcal{T}^{\pi}Q^{\pi}\right](s,a)-\widehat{Q}(s,a)+\left[\mathcal{T}^{\pi}\widehat{Q}\right](s,a)\right]$
$=\frac{1}{1-\gamma}\operatorname{\mathbb{E}}_{d^{\pi}}\left[\left[\mathcal{T}^{\pi}\widehat{Q}\right](s,a)-\widehat{Q}(s,a)\right]$
$\leq\frac{\mathcal{C}_{\infty}^{\pi}}{1-\gamma}\cdot\operatorname{\mathbb{E}}_{\mu}\left[\left\lvert\left[\mathcal{T}^{\pi}\widehat{Q}\right](s,a)-\widehat{Q}(s,a)\right\rvert\right$]
$\leq\max_{g\in\mathcal{G}_{\widehat{Q}}}\operatorname{\mathbb{E}}_{\mu}\left[\textrm{sgn}\left(\widehat{Q}(s,a)-g(s,a)\right)\left(\widehat{Q}(s,a)-r-\gamma\widehat{Q}(s^{\prime},\pi)\right)\right]$
@@
${\textrm{sgn}\left(Q(s,a)-g(s,a)\right)\left(Q(s,a)-r-\gamma Q(s^{\prime},\pi)\right)}\in\left[-V_{\max},V_{\max}\right],$
@@
$\Big|\operatorname{\mathbb{E}}_{\mu}\left[\textrm{sgn}\left(\widehat{Q}(s,a)-g(s,a)\right)\left(\widehat{Q}(s,a)-r-\gamma\widehat{Q}(s^{\prime},\pi)\right)\right]$
$~-\operatorname{\mathbb{E}}_{\mathcal{D}}\left[\textrm{sgn}\left(\widehat{Q}(s,a)-g(s,a)\right)\left(\widehat{Q}(s,a)-r-\gamma\widehat{Q}(s^{\prime},\pi)\right)\right]\Big|\leq 2V_{\max}\sqrt{\frac{\log\left(\frac{2m}{\delta}\right)}{n}}\vcentcolon={}\varepsilon_{\mathsf{obj}}.$
@@
$J(\pi)-\operatorname{\mathbb{E}}_{d_{0}}\left[\widehat{Q}(s,\pi)\right]$ $\leq\max_{g\in\mathcal{G}_{\widehat{Q}}}{}\operatorname{\mathbb{E}}_{\mathcal{D}}\left[\textrm{sgn}\left(\widehat{Q}(s,a)-g(s,a)\right)\left(\widehat{Q}(s,a)-r-\gamma\widehat{Q}(s^{\prime},\pi)\right)\right]+\varepsilon_{\mathsf{obj}}$
$\leq\max_{g\in\mathcal{G}_{Q^{\pi}}}{}\operatorname{\mathbb{E}}_{\mathcal{D}}\left[\textrm{sgn}\left(Q^{\pi}(s,a)-g(s,a)\right)\left(Q^{\pi}(s,a)-r-\gamma Q^{\pi}(s^{\prime},\pi)\right)\right]+\varepsilon_{\mathsf{obj}}$
$\leq\max_{g\in\mathcal{G}_{Q^{\pi}}}{}\operatorname{\mathbb{E}}_{\mu}\left[\textrm{sgn}\left(Q^{\pi}(s,a)-g(s,a)\right)\left(Q^{\pi}(s,a)-r-\gamma Q^{\pi}(s^{\prime},\pi)\right)\right]+2\cdot\varepsilon_{\mathsf{obj}}$
$=\max_{g\in\mathcal{G}_{Q^{\pi}}}{}\operatorname{\mathbb{E}}_{\mu}\left[\textrm{sgn}\left(Q^{\pi}(s,a)-g(s,a)\right)\left(Q^{\pi}(s,a)-\left[\mathcal{T}^{\pi}Q^{\pi}\right](s,a)\right)\right]+2\cdot\varepsilon_{\mathsf{obj}}$
$=\max_{g\in\mathcal{G}_{Q^{\pi}}}{}\operatorname{\mathbb{E}}_{\mu}\left[\textrm{sgn}\left(Q^{\pi}(s,a)-g(s,a)\right)\left(Q^{\pi}(s,a)-Q^{\pi}(s,a)\right)\right]+2\cdot\varepsilon_{\mathsf{obj}}$
$=2\cdot\varepsilon_{\mathsf{obj}},$
@@
$\widehat{Q}=\operatorname*{arg\,min}_{Q\in\mathcal{Q}}\max_{g\in\mathcal{G}_{Q}}\operatorname{\mathbb{E}}_{\mathcal{D}}\left[\textrm{sgn}\left(Q(s,a)-g(s,a)\right)\left(Q(s,a)-r-Q(s^{\prime},\pi)\right)\right].$
@@
$J(\pi)-\operatorname{\mathbb{E}}_{d_{0}}\left[\widehat{Q}(s,\pi)\right]\leq 2\cdot\mathcal{C}_{\infty}^{\pi}\cdot\varepsilon_{\mathsf{obj}}.$
@@ C.5 Further Discussion on Comparison to BVFT and the Coverage Assumptions
As mentioned above, the coverage condition of LSTD-Tournament is inherited from its “base” algorithm, i.e., the $\sigma_{\min}(A)$ term in LSTDQ. Despite its wide use in the analyses of LSTDQ algorithms [30, 3, 48], we discover several problems with this definition that warrants further investigation. First, the quantity is not invariant to reparameterization, in the sense that if we scale the linear features $\phi$ with a constant (and scale $\theta$ accordingly), $\sigma_{\min}(A)$ will also change despite that the estimation problem essentially remains the same. Second, if $\phi$ has linearly dependent coordinates, $\sigma_{\min}(A)$ can be 0 while this says nothing about the quality of data (which could have perfect coverage); this is usually not a concern in LSTDQ analysis since we can always pre-process the features, but for its application in LSTD-Tournament the lack of linear independence can happen if we have nearly identical candidate functions $Q_{i}\approx Q_{j}$. Last but not least, we lack good understanding of when $\sigma_{\min}(A)$ is well-behaved outside very special cases; for example, it is well-known that fully on-policy data (in the sense that $\mu$ is invariant to the transition dynamics under $\pi$) implies lower-bounded $\sigma_{\min}(A)$. However, as soon as $\mu$ is off-policy, there lacks general characterization of what kind of $\mu$ helps lower-bound $\sigma_{\min}(A)$. For example, when concentrability coefficient $\mathcal{C}_{\infty}^{\pi}$ is well bounded, it is unclear whether that implies lower-bounded $\sigma_{\min}(A)$.
@@ D.2 Experiment Settings
In the above setup, the behavior and the target policies all stem from the same DDPG training procedure. While these policies still have significant differences (see Figure 2L), the distribution shift is relatively mild. For the data coverage experiments (Section 6.3), we prepare a different set of behavior policies that intentionally offer poor coverage: these policies, denoted as $\pi_{i}^{\textrm{poor}}$, are obtained by running DDPG with a different neural architecture (than the one used for generating $\pi_{0:14}$) in a different environment of $g=-60,\sigma=100$. We also provide the parallel of our main experiments in Figures 3 and 4 under these behavior policies with poor coverage in Appendix E.1.
@@
Gravity $\mathbf{g}$ Noise Level $\sigma$
Groundtruth Model $M^{\star}$ and
Behavior Policy $\pi_{b}$
Target
Policies $\Pi$